A proton makes a (radius = 30.0 cm) circular orbit in a 2.5 T magnetic field. What is the velocity of the proton?

mass of proton = 1.67x10-27 kge = 1.602x10-19 C

1
Expert's answer
2022-05-12T14:29:03-0400
"\frac{mv^2}{r}=qvB"

"v=\frac{qBr}{m}=\frac{1.6*10^{-19}*2.5*0.30}{1.67*10^{-27}}=7.2*10^7\:\rm m/s"